OpenStudy (anonymous):

a table for y=2x^2-8 is given solve each equation. a. 2x^2-8=0,b. 2x^2-8<0,c. 2x^2-8>0.

OpenStudy (anonymous):

where is the table?

OpenStudy (anonymous):

it is in the question

OpenStudy (anonymous):

2x^2-8=0 2x^2=8 x^2=4 x=-2 or 2

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

is that forthe equation a.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what about b and c

OpenStudy (anonymous):

Did u understand it?

OpenStudy (anonymous):

yes for thef irst one I did

OpenStudy (anonymous):

2x^2-8<0 2x^2<8 x^2<4 x^2<(+-2)^2 Now, IF x^2<(+-a)^2 then -a<x<a

OpenStudy (anonymous):

SO, can u complete it?

OpenStudy (anonymous):

no do not get that one

OpenStudy (anonymous):

2x^2-8<0 2x^2<8 x^2<4 x^2<(+-2)^2 IS it OKAY up to here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would it be 4

OpenStudy (anonymous):

IF x^2<(+-a)^2 then -a<x<a

OpenStudy (anonymous):

try to use it

OpenStudy (anonymous):

I do notg et it

OpenStudy (anonymous):

x^2<(+-2)^2 -2<x<2

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes but how do i figure it out

OpenStudy (anonymous):

HINT: IF x^2<(+-a)^2 then -a<x<a

OpenStudy (anonymous):

It is like formula

OpenStudy (anonymous):

howdo i figurethef o rmula out

OpenStudy (anonymous):

U want me to derive it?

OpenStudy (anonymous):

what does that mean

OpenStudy (anonymous):

hello what doyou m ean

OpenStudy (anonymous):

I am sorry but I guess u need to revise your algebra once

OpenStudy (anonymous):

soyoucannoth elp meany further

OpenStudy (anonymous):

Sorry not so good at explaining things

OpenStudy (anonymous):

ok Iwil l try someelse

OpenStudy (anonymous):

@mathslover can help u?

OpenStudy (anonymous):

He is good at explaining

mathslover (mathslover):

I am well at explanation :)

OpenStudy (anonymous):

socanyouhelpme mathsolver

mathslover (mathslover):

I will be right back ... wait for 2 minutes please

OpenStudy (anonymous):

\[ 2x^2-8<0 \\ 2x^2<8 \\ x^2<4 \\ \sqrt{x^2}<\sqrt{4} \]Now it is important to remember that \(\sqrt{\quad}\) mean the "positive square root" (not the negative one) and that \(\sqrt{a^2}\) is the definition of of the absolute value.\[ |x| < 2 \]

OpenStudy (anonymous):

To solve absolute value equations, we have to split it up into two cases. Case 1: Assume \(x\) (the thing in the absolute value) is positive. \[ |x| < 2 \\ x<2 \] That was easy. Case 2: Assume \(x\) is negative: \[ |x| < 2 \\ -x < 2 \\ x > -2 \]Remember that if \(x\) is negative, then \(|x| =-x\) because the absolute value bars had to flip the sign to make \(x\) positive. Also remember that when you multiply/divide both sides of an inequality by a negative number, the equality flips. Hence \(<\) became \(>\) in this instance.

OpenStudy (anonymous):

Notation allows us to conveniently combine the expressions \(x < 2\) and \(x > -2\) into \(-2 < x < 2\). Does this help, @cdelomas ?

OpenStudy (anonymous):

One thing worthy of committing to memory is that: \[ |x| < a \implies -a < x < a \\ |x| > a \implies x < -a,\quad x > a \]

OpenStudy (anonymous):

so for equation b is it squareroot x2>squareroot 4.